OpenStudy (anonymous):

What are the x– and y–intercepts of the quadratic equation y = x2 + 2x – 3?

OpenStudy (anonymous):

Set y=0 and solve for x to find x-intercept and set x=0 to find y-intercept

OpenStudy (anonymous):

So, 0=x^2+2x-3 0=(x-1)(x+3) Either x=1 or x=-1 y=-3

OpenStudy (anonymous):

For the y-intercept, plug in 0 for x. Y-Intercept = -3 Then, plug in 0 for y and solve the quadratic equation. 0 = x^2 + 2x - 3 0 = (x + 3)(x - 1) x + 3 = 0 x = -3 x - 1 = 0 x = 1 Therefore, the x-intercepts are -3 and 1.